OpenStudy (anonymous):

for the equation -5(2x+2) greater than or equal to -10x-17 is the solution x is greater than or equal to 7, x is less than or equal to 19, all real numbers or no solution

OpenStudy (anonymous):

\[-5(2x+2) \geq -10x-17\] \[-10x -10 \geq -10x - 17\] \[-10\geq -17\] which is valid. The point being, the -10x on either side 'cancel' and so for all intents and purposes, the inequality is true for all real numbers. If it had no solution, the inequality would somehow have to be violated which is not the case...

OpenStudy (anonymous):

thank you soooo much! that helped a lot